3(2x + 11) = 9 Reason for each step

6x + 33 = 9 Reason for each step

6x=-24 reason for each step

X=-4 reason for each step


44 - 2(3x + 4) = -18x Reason for each step


44 - 6x - 8 = -18x Reason for each step

-6x + 36 = -18x Reason for each step

36 = -12x Reason for each step


-3 = x Reason for each step
1
Expert's answer
2020-09-28T09:19:35-0400

1.

Simplifying

"3(2x + 11) = 9"


Reorder the terms:

"3(11 + 2x) = 9"

"(11 * 3 + 2x * 3) = 9"

"(33 + 6x) = 9"


Solving

"33 + 6x = 9"


Solving for variable 'x'.

Move all terms containing x to the left, all other terms to the right.


Add '-33' to each side of the equation.

"33 +( -33) + 6x = 9 + (-33)"


Combine like terms: "33 + (-33) = 0"

"0 + 6x = 9 +(-33)"

"6x = 9 +( -33)"


Combine like terms: "9 + (-33) = -24"

Divide each side by '6'.

"6x = -24"


Simplifying

"x = -4"


2.

Simplifying

"44 + (-2)(3x + 4) = -18x"


Reorder the terms:

"44 + (-2)(4 + 3x) = -18x"

"44 + (4 *( -2) + 3x *(-2)) = -18x"

"44 + (-8 + (-6x)) = -18x"


Combine like terms: "44 +( -8) = 36"

"36 +( -6)x = -18x"


Solving

"36 +( -6)x = -18x"


Solving for variable 'x'.


Move all terms containing x to the left, all other terms to the right.


Add '18x' to each side of the equation.

"36 + (-6)x + 18x = -18x + 18x"


Combine like terms:"-6x + 18x = 12x"

"36 + 12x = -18x + 18x"


Combine like terms: "-18x + 18x = 0"

"36 + 12x = 0"


Add '-36' to each side of the equation.

"36 + (-36) + 12x = 0 +(-36)"


Combine like terms: "36 + (-36) = 0"

"0 + 12x = 0 +(-36)"

"12x = 0 + (-36)"


Combine like terms:"0 + (-36) = -36"

"12x = -36"


Divide each side by '12'.

"x = -3"


Simplifying

"x = -3"